OpenStudy (anonymous):

8/5w-10=3/5w how do i solve this?

OpenStudy (anonymous):

w=-1/10

OpenStudy (anonymous):

can you show how you got the anwser plz?

OpenStudy (anonymous):

-10=3/(5w)-8/(5w) -10=-5/(5w) -10=-1/w -10w=-1 w=1/10

OpenStudy (anonymous):

im kinda confused? lol

OpenStudy (anonymous):

first take the terms with unknown to one side and the once with just number to the other

OpenStudy (anonymous):

then combine like term

OpenStudy (anonymous):

8/5 & 3/5 are fractions.... just making sure cuz ppl get confused.

OpenStudy (anonymous):

yes I got that ok then now tell me where is the w is it in denominator or numerator cauze this solution is for w in denominator

OpenStudy (anonymous):